OpenStudy (anonymous):

Find the length of Acr AB: if the angle measure is 65 degrees and the diameter is 14in

OpenStudy (anonymous):

\[\frac{AB}{Circumference}=\frac{65^o}{360^o} \rightarrow \frac{65 \times 2\pi(14/2)}{360} \approx 7.94 inches\]